OpenStudy (stormoshawty):

What is the solution of the system? y=14x-1 y=-11x+4

OpenStudy (anonymous):

multiply the 1st equation with -1 -y=-14x+1 y=-11x+4 add equations 0=-25x+5 25x=5, x=0.2

OpenStudy (anonymous):

and plug 0.2 into one of the equations to obtain y.

OpenStudy (stormoshawty):

The choices are \[(5, 51) \] \[(\frac{ 1 }{ 5 }, \frac{ 9 }{ 5 })\] \[(\frac{ 5 }{ 3 }, \frac{ 67 }{ 3 })\] \[(\frac{ 3 }{ 5 }, \frac{ 37 }{ 5 })\]

OpenStudy (unklerhaukus):

to get the right option in your question use the same method strenesmee has suggested, but use fractions rather than decimals